# Local wind speed forecasting at short time horizons relying on both Numerical Weather Prediction and observations from surrounding stations

### v1

@@ 3.1 Prediction at multiple horizons
Let $V^{s}_{d,t}$ be the surface wind velocity (in $m\;s^{-1}$) at time $t$ of day $d$ at a given fixed location $s$. In all our experiments we consider the hourly frequency and thus, for a given day $d$, $t$ is expressed in hours. The upper character $s$ can represent the longitude/latitude coordinates of the target location or simply the site name. Notice that $V^{s}_{d,t}$ is directly recorded if $s$ corresponds to a site where a ground station is installed. Let $h_{1}<h_{2}<\ldots<h_{H}$ be a fixed set of $H$ increasing horizon values with $H\geq 1$. For the sake of simplicity and to avoid handling horizons too far away from NWP run initial time, we restrict ourselves to the set of horizon times $h\in\{1,2,3,4,5,6\}$ hours. This corresponds to $H=6$ with $\{h_{i}=i\}_{i=1,\ldots,6}$ in the previous notations.
We denote by ${\bf{V}}^{s}_{d,t+h}$ the $H$-dimensional vector of future velocities at multiple horizons:
${\bf{V}}^{s}_{d,t+h}=\begin{pmatrix}V^{s}_{d,t+h_{1}}\\ V^{s}_{d,t+h_{2}}\\ \vdots\\ V^{s}_{d,t+h_{H}}\end{pmatrix}\;.$
Let us notice that the horizon time $h$ can be also called forecast horizon or simply the horizon. In the literature $h$ is also often referred to as the lead time (e.g. in Marcille et al. (2024)). The time $t$ of the day when the prediction is performed will be called the initial time. The objective of all models discussed below is, for any given station $s$, any day $d$ and initial time $t$, to provide the “best” prediction of upcoming velocity vector ${\bf{V}}^{s}_{d,t+h}$.
@@
It is important to realize that, since in the MeteoNet dataset we only have access to NWP models initialized at 00 UTC (hereafter referred to as the initialization time), for a given initial time $t$ and forecast horizon $h$, their predictions correspond to an effective horizon of $t+h$, which places them at a disadvantage. The experiment we conduct does not aim to compare NWPs in general—since these may be re-initialized multiple times per day—but rather those specifically from the 00 UTC run. For obvious reasons, we choose to compare results at a fixed forecast horizon $h$. Indeed, from an operational perspective, for any given initial $t$ time within the day, this is the most relevant variable. Consequently, the conclusions of this study are based on the average performance over all $t$ at fixed $h$, rather than at fixed $t+h$, thus encompass weather phenomena that occur at different times of the day. In this framework, the performance of models that involve only AROME or ARPEGE predictions effectively integrates a range of effective horizons, which depends only marginally on $h$. As a result, we expect NWP-based results to exhibit very limited dependence on $h$.
The prediction vector will be denoted as ${\widehat{{\bf{V}}}}^{s}_{d,t+h}$. In general, at each day $d$ and time $t$, the “best” predictor corresponds to the one that minimizes the conditional Mean Squared Error (MSE):
$\mathcal{E}^{mse}_{d,t}={\mathbb{E}}\left(\left\|{\bf{V}}^{s}_{d,t+h}-{\widehat{{\bf{V}}}}^{s}_{d,t+h}\right\|^{2}\Big{|}\;{\cal F}_{d,t}\right)$ (1)
where ${\cal F}_{d,t}$ represents all the “information” available at day $d$, time $t$ and $\left\|{\bf{Z}}\right\|^{2}$ is the squared euclidian norm of ${\bf{Z}}$, i.e., the sum of its squared components: $\sum_{i=1}^{H}Z_{i}^{2}$. $\mathbb{E}(.|{\cal F}_{d,t})$ stands for the mathematical expectation conditionnally to ${\cal F}_{d,t}$ meaning the expectation associated with some (unobservable) conditional probability law of future outcome of weather conditions at times $t+h$ that could be, for instance, defined within a Bayesian approach.
@@
It can be shown that the vector ${\bf{M}}^{s}_{d,t+h}$ that minimizes $\mathcal{E}^{mse}_{d,t}$ is simply the conditional mean:
${\bf{M}}^{s}_{d,t+h}=\mathbb{E}\left({\bf{V}}^{s}_{t+h}\Big{|}\;{\cal F}_{d,t}\right)\;.$
@@
In that respect, for any station $s$, at any time $t$ and day $d$, minimizing the MSE is equivalent to building ${\widehat{{\bf{V}}}}^{s}_{d,t+h}$ as the best approximation of the conditional expectation ${\bf{M}}^{s}_{d,t+h}$ of the multi-horizon velocity vector ${\widehat{{\bf{V}}}}^{s}_{d,t+h}$.
@@
In practice, at each day $d$ and time $t$, one “encodes” the information ${\cal F}_{d,t}$ available at site $s$ within a large vector of observable “features” ${\bf{X}}^{s}_{d,t}$. A commonly used possibility is then to seek for a model ${\mathcal{M}}({\bf{X}}^{s}_{d,t},{\bf{\Theta}})$ that directly outputs ${\bf{V}}^{s}_{d,t+h}$:
${\widehat{{\bf{V}}}}^{s}_{d,t+h}={\mathcal{M}}({\bf{X}}^{s}_{d,t},{\bf{\Theta}})$ (2)
where ${\bf{\Theta}}$ is a set of model parameters.
Another option is to operate within the framework of the so-called “probabilistic forecasting”. Instead of providing a direct estimation of the conditional mean, the model provides ${\widehat{\rho}}^{s}_{d,t+h}({\bf{V}})$, an estimation of the conditional law of ${\bf{V}}_{d,t+h}^{s}$. As explained below, this can be done by various approaches like e.g. a direct output of a set of quantiles values of ${\widehat{\rho}}^{s}_{d,t+h}$ or within a parametric method. In that case, the conditional mean ${\widehat{{\bf{V}}}}_{t+h}^{s}$ is obtained as follows:
${\widehat{\rho}}^{s}_{d,t+h}$ $=$ ${\mathcal{M}}({\bf{X}}^{s}_{d,t},{\bf{\Theta}})\;,$ (3)
${\widehat{{\bf{V}}}}^{s}_{d,t+h}$ $=$ $\int_{{\mathbb{R}}^{H}}{\widehat{\rho}}^{s}_{d,t+h}({\bf{V}})\;.{\bf{V}}\;dV_{1}\ldots dV_{H}$ (4)
In the following sections we introduce a specific model ${\mathcal{M}}$ that is built using a neural network architecture and whose final layer is configured to match the context described by Eq. (2) and a similar model ${\mathcal{M}}^{\prime}$ such that the output layer corresponds to the framework of Eq. (3). We also consider some of its simple variants together with baseline models for comparison purpose.
3.2 A neural network model designed to handle station data and NWP predictions
Figure 2: Sketch of neural network architecture we use to handle station, AROME and ARPEGE data. Output shape is indicated under each layer. F.C. stands for “fully connected layer” and $LSTM^{*}$ means that the full sequence of outputs is returned and not only the final one.
The objective of this study is to explore the hybridization of NWP and observations using a streamlined neural network model. Rather than developing a state-of-the-art model, we employ a straightforward architecture with standard convolutional and recurrent layers for spatial and temporal feature extraction (Goodfellow et al., 2016). Identifying advanced architectures and optimizing model performance is left for future research.

### v2

@@ 3 Prediction task at multiple horizons
Let $V^{s}_{d,t}$ be the surface wind velocity (in $m\;s^{-1}$) at time $t$ of day $d$ at a given fixed location $s$. In all our experiments we consider the hourly frequency and thus, for a given day $d$, $t$ is expressed in hours. The upper character $s$ can represent the longitude/latitude coordinates of the target location or simply the site name. Notice that $V^{s}_{d,t}$ is directly recorded if $s$ corresponds to a site where a ground station is installed. The time $t$ of the day when the prediction is performed will be called the initial time. Let ${\mathbf{H}}=(h_{1},h_{2},\ldots,h_{N_{H}})$ be a set of $N_{H}$ fixed time horizons values such that $h_{1}<h_{2}<\ldots<h_{N_{H}}$. For the sake of simplicity and to avoid handling horizons too far away from NWP run initial time, we restrict ourselves to the set of next $N_{H}=6$ upcoming hours, namely ${\mathbf{H}}=(1,2,3,4,5,6)$ hours. We denote by ${\mathbf{V}}^{s}_{d,t+{\mathbf{H}}}$ the $6$-dimensional vector of future velocities at multiple horizons:
${\mathbf{V}}^{s}_{d,t+{\mathbf{H}}}=\begin{pmatrix}V^{s}_{d,t+1}\\ V^{s}_{d,t+2}\\ \vdots\\ V^{s}_{d,t+6}\end{pmatrix}\;$.
In the literature a single horizon time $h_{i}$ is also often referred to as the lead time (e.g. in Marcille et al. (2024)). The forecasting task we consider is therefore to compute, at each $d$ and $t$, for each horizon $h_{i}\in{\mathbf{H}}$, a velocity value $\widehat{V}^{s}_{d,t+h_{i}}$ that estimates “at best” $V^{s}_{d,t+h_{i}}$. The corresponding 6-dimensional prediction vector will be denoted as ${\widehat{{\mathbf{V}}}}^{s}_{d,t+{\mathbf{H}}}$. Notice that for a single horizon $h$ and initial time $t$, the time $t+h$ is often referred to as the valid time or the target time.
@@
At this stage, it is important to mention that our goal is not to compare models relying on NWP in general, which may be initialized multiple times per day, but specifically those from the 00 UTC run. Indeed, the MeteoNet dataset only includes NWP forecasts initialized at 00 UTC (hereafter referred to as the initialization time), a forecast made at time $t$ with horizon $h$ corresponds therfore to an “effective horizon” of $t+h$. In the following, we evaluate performance at fixed forecast horizons $h$, which is more relevant operationally than fixing the effective time $t+h$. This means we average results over all initial times $t$, at fixed $h$, thereby covering weather events occurring at different times of day. In this framework, the performance of models that involve only AROME or ARPEGE predictions effectively integrates a range of effective horizons, which depends only marginally on $h$ for t large enough. As a result, we expect NWP-based results to exhibit very limited dependence on $h$.
The most commonly used definition of the “best prediction” ${\widehat{{\mathbf{V}}}}^{s}_{d,t+{\mathbf{H}}}$ is the one that minimizes the Mean Squared Error (MSE):
$\mathcal{E}^{mse}\!\!={\mathbb{E}}\left(\left\|{\mathbf{V}}^{s}_{d,t+{\mathbf{H}}}-{\widehat{{\mathbf{V}}}}^{s}_{d,t+{\mathbf{H}}}\right\|^{2}\right)=\sum_{i=1}^{6}{\mathbb{E}}\left(\left|{\widehat{V}}^{s}_{d,t+h_{i}}-V^{s}_{d,t+h_{i}}\right|^{2}\right)$ (1)
where $\mathbb{E}$ represents the mathematical expectation and $\left\|{\mathbf{Z}}\right\|^{2}$ is the squared euclidian norm of $\mathbf{Z}$}. Along the same line, one defines the conditional mean squared error as:
$\mathcal{E}^{mse}_{d,t}$ $=$ ${\mathbb{E}}\left(\left\|{\mathbf{V}}^{s}_{d,t+{\mathbf{H}}}-{\widehat{{\mathbf{V}}}}^{s}_{d,t+{\mathbf{H}}}\right\|^{2}\Big|\;{\cal F}_{d,t}\right)$ (2)
$=$ $\sum_{i=1}^{6}{\mathbb{E}}\left(\left|{\widehat{V}}^{s}_{d,t+h_{i}}-V^{s}_{d,t+h_{i}}\right|^{2}\Big|\;{\cal F}_{d,t}\right)$
where ${\cal F}_{d,t}$ represents all the “information” available at day $d$, time $t$ and $\mathbb{E}(.|{\cal F}_{d,t})$ stands for the mathematical expectation conditionnally to ${\cal F}_{d,t}$ meaning the expectation as respect to the conditional probability law of future outcome of weather conditions at times $t+h_{i}$, $h_{i}\in{\mathbf{H}}$. In practice, at each day $d$ and time $t$, the abstract information set $\mathcal{F}_{d,t}$ available at site $s$ is encoded into a high-dimensional feature vector ${\mathbf{X}}^{s}_{d,t}$. This vector may include, for instance, weather observations at station $s$ and its neighboring stations, the day of the year, and other relevant variables (see next section for details). Accordingly, the conditional expectation $\mathbb{E}(\cdot\mid\mathcal{F}_{d,t})$ is approximated by $\mathbb{E}(\cdot\mid{\mathbf{X}}^{s}_{d,t})$.
From the law of iterated expectations, we simply have
$\mathcal{E}^{mse}={\mathbb{E}}\left(\mathcal{E}^{mse}_{d,t}\right)$
and therefore minimizing at all $t,d$, $\mathcal{E}^{mse}_{d,t}$ leads to minimizing the global error $\mathcal{E}^{mse}$.
@@
It is well known that the vector ${\mathbf{M}}^{s}_{d,t+{\mathbf{H}}}$ that minimizes $\mathcal{E}^{mse}_{d,t}$ is simply the conditional mean:
${\mathbf{M}}^{s}_{d,t+{\mathbf{H}}}=\mathbb{E}\left({\mathbf{V}}^{s}_{t+{\mathbf{H}}}\Big|\;{\mathbf{X}}^{s}_{d,t}\right)\;.$
@@
In that respect, for any station $s$, at any time $t$ and day $d$, minimizing the MSE is equivalent to building ${\widehat{{\mathbf{V}}}}^{s}_{d,t+{\mathbf{H}}}$ as the best approximation of the conditional expectation ${\mathbf{M}}^{s}_{d,t+{\mathbf{H}}}$ of the multi-horizon velocity vector ${\widehat{{\mathbf{V}}}}^{s}_{d,t+{\mathbf{H}}}$.
As mentioned above, we will consider two distinct methods to do that, a deterministic approach and a probalistic appraoch.
3.1 Deterministic approach
@@
A commonly used possibility to estimate ${\mathbf{M}}^{s}_{d,t+{\mathbf{H}}}$ is to seek for a model ${\mathcal{M}}({\mathbf{X}}^{s}_{d,t},{\mathbf{\Theta}})$ that directly outputs ${\mathbf{V}}^{s}_{d,t+h}$:
${\widehat{{\mathbf{V}}}}^{s}_{d,t+{\mathbf{H}}}={\mathcal{M}}({\mathbf{X}}^{s}_{d,t},{\mathbf{\Theta}})$ (3)
where $\mathbf{\Theta}$} is a set of model parameters which is chosen in order to minimize $\mathcal{E}^{mse}$. This latter can be estimated empirically as the following loss function:
$\mathcal{L}^{mse}$ $=$ $N^{-1}\sum_{k=1}^{N}\left\|{\mathbf{V}}^{s}_{d,t_{k}+{\mathbf{H}}}-{\widehat{{\mathbf{V}}}}^{s}_{d,t_{k}+{\mathbf{H}}}\right\|^{2}$ (4)
$=$ $N^{-1}\sum_{k=1}^{N}\sum_{i=1}^{6}\left({\widehat{V}}^{s_{k}}_{d_{k},t_{k}+h_{i}}-V^{s_{k}}_{d_{k},t_{k}+h_{i}}\right)^{2}$
where $N$ is the number of available samples (in the training subset, see below) which corresponds to a fixed fraction of the total number of different triplets “station”, “day” and “initial time” ($s$,$d$,$t$) for which the observed velocity vector ${\mathbf{V}}^{s}_{d,t+{\mathbf{H}}}$ and the input features ${\mathbf{X}}^{s}_{d,t}$ are available.
To assess the forecasting performance of a model, the most commonly used score is the Root Mean Square Error (RMSE), defined as:
$\text{RMSE}=\sqrt{\mathcal{L}^{mse}}$ (5)
where $\mathcal{L}^{mse}$ is computed on either the validation or test subset. RMSE corresponds to the empirical deviation (in $m\;s^{-1}$) of observed velocity as respect to the prediction $^{\widehat{{\mathbf{V}}}}$.
3.2 Probabilistic approach
Alternatively to the previous deterministic (also called “single-point”) forecast, one can estimate the full conditional probability distribution from which the conditional mean is then derived (see Eqs. (6), (8)). Probabilistic learning methods has become increasingly important in recent years as it captures not only the expected outcome but also uncertainties around this expectation and therefore provides notably insights into the likelihood of extreme events. They can rely on different approaches, like quantile-based methods which consists in directly estimating specific quantiles as the median or the 90th percentile, without assuming any particular probability distribution (see, e.g., Koenker (2005); Park et al. (2022)). Another widely used strategy involves coupling a sequential model with a likelihood model, where the latter governs how observations are generated from underlying hidden or latent states (see, e.g., Rasmussen and Williams (2005); Rangapuram et al. (2018)). The method we use in this work relies on such approach and is inspired by Salinas et al. (2020). It consists in estimating, at each time step, the set of parameters characterizing the conditional law of upcoming velocity chosen to belong to a given parametric class of distributions. This approach is very parsimonious, offering a complete characterization of the forecast distribution, and it can easily integrate additional information, such as prior distribution, using within the framework of Bayesian inference. More specifically, for each station $s$, day $d$ and initial time step $t$, the tensor of parameters ${\mathbf{\Pi}}_{s,d,t}$ characterizing, $\rho({\mathbf{V}};{\mathbf{\Pi}}_{s,d,t})$, the conditional law of ${\mathbf{V}}^{s}_{d,t+{\mathbf{H}}}$ as the output of a neural-network model $\mathcal{M}$$^{\prime}$:
${\mathbf{\Pi}}_{s,d,t}={\mathcal{M}}^{\prime}({\mathbf{X}}^{s}_{d,t},{\mathbf{\Theta}})\;$ (6)
where again, $\mathbf{\Theta}$} stands for the set of parameters characterizing the model $\mathcal{M}$$^{\prime}$ that have to be learned. For that purpose we suppose that the multivariate density is simply the product of monovariate densities:
$\rho({\mathbf{V}};{\mathbf{\Pi}}_{s,d,t})=\prod_{i=1}^{6}\rho(V_{i};{\mathbf{\Pi}}^{i}_{s,d,t})$
where ${\mathbf{\Pi}}^{i}_{s,d,t}$ is the vector parameters of the conditional law of $V^{s}_{d,t+h_{i}}$. The model $\mathcal{M}$$^{\prime}$ parameters are learnt by minimizing the negative log-likelihood (also called logarithmic score):
$\mathcal{L}^{\ell}=-N^{-1}\sum_{k=1}^{N}\sum_{i=1}^{6}\ln\rho(V^{s_{k}}_{d_{k},t_{k}+h_{i}};{\mathbf{\Pi}}^{i}_{t_{k}})\;.$ (7)
In this framework, the conditional mean estimation ${\widehat{{\mathbf{V}}}}_{t+{\mathbf{H}}}^{s}$ is computed as follows:
${\widehat{V}}^{s}_{d,t+h_{i}}=\int_{{\mathbb{R}}}z\rho(z;{\mathbf{\Pi}}^{i}_{s,d,t})\;dz\;.$ (8)
We choose to base our computations on $\rho(z)$ belonging to the class of so-called “Multifractal-Rice” (M-Rice) distribution, which has proven to be particularly suitable for wind speed forecasting when compared to other distributions like Weibull or Gamma (Baggio and Muzy, 2024). The M-Rice distribution, introduced in Baïle et al. (2011), is motivated by the random cascade framework used in fully developed turbulence. It extends the classical Rice distribution by allowing its scale parameter to be random (typically modeled as a log-normal variable). The resulting distribution is governed by three parameters: $(\nu,\sigma^{2},\lambda^{2})$. Here, $\nu$ represents a mean parameter, $\sigma^{2}$ a dispersion parameter, and $\lambda^{2}$, often referred to as the intermittency parameter, controls the shape of the distribution. In particular, larger values of $\lambda^{2}$ lead to heavier tails, capturing the increased probability of extreme values. The mathematical definition of M-Rice distribution, the interpretation of its parameters and details for numerical estimations are provided in A.
3.3 Extreme event prediction and associated performance scores
As already emphasized, since our goal is to compare models providing single-point predictions, we treat model $\mathcal{M}$$^{\prime}$ of previous section that provides a probabilistic forecasting as an intermediate step for generating deterministic predictions for example through Eq. (8) which performance can be evaluated using the RMSE metrics defined in (5).
[…]
$I_{s,d,t}(Q)={\mathcal{H}}(V^{s}_{d,t}-Q)=\begin{cases}1&\text{if }V^{s}_{d,t}\geq Q\\ 0&\text{otherwise}.\end{cases}$ (9)
Within this context, predicting strong wind events at a given location amounts to forecasting the occurrence of threshold exceedances with some threshold corresponding to a high quantile relative to the site’s wind climatology. Let $Q_{C}^{s}(p)$ denote such a quantile of probability level $p$ close to $p=1$. If the climatological distribution at station $s$ is denoted as $\rho^{s}_{C}(V)$, one has, by definition:
$\int_{0}^{Q_{C}^{s}(p)}\rho^{s}_{C}(V)\;dV=p\;.$ (10)
In that respect, if $1-p\ll 1$, $I_{s,d,t}(Q_{C}^{s}(p))$ indicates the occurrence of strong wind speeds greater than a threshold in the tail of the site’s climatological distribution (i.e., corresponding to a high level quantile). Predicting strong events occurrence thus amounts to predict the value of $I_{s,d,t+h}(Q^{s}_{C}(p))$ at each time $t$, day $d$, for each station $s$ and horizon $h$.
Assessing the quality of forecasts in this context thus amounts to evaluating the accuracy of binary predictions which is a classification task—in contrast to the regression setting considered earlier. In this classification framework, various skill scores can be used to evaluate performance, depending on the aspect of predictive quality one wishes to emphasize (e.g., sensitivity, specificity, overall accuracy). For a comprehensive overview of relevant metrics and their properties, we refer the reader to Jolliffe (2004) and Wilks (2019). Among all possible scores, the Peirce Skill Score (PSS) is a widely used metric for evaluating the performance of a binary classification model, particularly in the context of severe weather prediction. It is defined as:
$\text{PSS}=\text{HR}-\text{FA}$ (11)
where HR stands for “Hit Rate” and FA for “False Alarm Rate”. These rates are derived from the contingency table of predictions, where the $N$ forecast/observation pairs $({\widehat{I}}^{s}_{d,t},I^{s}_{d,t})$ are divided in true positives TP, false positives FP, false negatives FN and true negatives TN (with TP+FP+FN+TN=N). More specifically, $\text{HR}=\frac{\text{TP}}{\text{TP}+\text{FN}}$ corresponds to the fraction of correctly predicted positive events, while $\text{FA}=\frac{\text{FP}}{\text{TN}+\text{FP}}$ is the fraction of wrongly predicted, i.e. predicted as positive, negative events. Unlike accuracy-based scores, PSS accounts for both missed events (false negatives) and false alarms (false positives), providing a balanced assessment of quantifying model performance. Moreover, this score is particularly suitable for forecasting extremes as it weights more strongly correct yes forecasts of rare events, thus discouraging artificial distortions towards the more common class. The PSS ranges from 1-1 to $1$, with $\text{PSS}=1$ corresponding to a perfect forecast and $\text{PSS}=0$ corresponding to unskilled, i.e., purely random, forecast. This means that a negative PSS value reflects prediction performance that is worse-than-random.
Another score commonly used when forecasting rare events (Jolliffe, 2004) is the Critical Success Index (CSI), also known as Threat Score. It is defined as:
$\text{CSI}=\frac{\text{TP}}{\text{TP}+\text{FP}+\text{FN}}\,$, (12)
corresponding to the ratio between the times $I^{s}_{d,t}=1$ has been forecasted correctly and the total number of times the event either happened, has been forecasted or both. The maximum value for CSI is 1, while the lowest possible value is 0. As TN are not considered in Eq. (12), this score is not artificially inflated by correct forecasts of the more common ”no” event. However, contrary to PSS, it is not an equitable score, which makes comparisons between binary events with different frequency of occurrence not straightforward. In the following, we refer mostly to the PSS when assessing performance, however, we will show some results in terms of CSI as well.
In the case of a single point prediction at site $s$, day $d$ and time $t$, whatever the considered score, a natural predictor of $I_{s,d,t+h_{i}}(Q_{C}^{s}(p))$ is simply:
${\widehat{I}}^{1}_{s,d,t+h_{i}}={\mathcal{H}}\Big({\widehat{V}}^{s}_{d,t+h_{i}}-Q_{C}^{s}(p)\Big)\;$ (13)
where ${\widehat{V}}_{d,t+h_{i}}^{s}$ is i-th component of ${\widehat{{\mathbf{V}}}}_{d,t+{\mathbf{H}}}^{s}$, the output of the considered model ($\mathcal{M}$} model or the conditional mean associated with the law predicted by $\mathcal{M}$$^{\prime}$ a probabilistic model). On the other hand, for any station $s$, the probabilistic approach provides, for each day $d$, initial time $t$ and horizon $h_{i}$, an estimate of ${\widehat{\rho}}^{s}_{d,t+h_{i}}(z)$, the conditional probability distribution of $V_{t+h_{i}}^{s}$ from which one can infer the conditional level $p$ quantile, ${\widehat{Q}}^{s}_{t+h_{i}}(p)$ that satisfies:
$\int_{0}^{\widehat{Q}^{s}_{t+h_{i}}(p)}{\widehat{\rho}}^{s}_{d,t+h_{i}}(z)\;dz=p\;.$ (14)
In this case, as shown in Mason (1979) (see also Jolliffe (2004)), one can explicitely build the “best” predictor of $I^{s}_{d,t+h}$, in the sense it is the one that maximizes the PSS, as:
${\widehat{I}}^{2}_{s,d,t+h}={\mathcal{H}}\Big(\widehat{Q}^{s}_{t+h_{i}}(p)-Q_{C}^{s}(p)\Big)\;.$ (15)
Such a prediction is different from the previous one in Eq. (13) since the expected conditional mean value estimation ${\widehat{V}}^{s}_{t+h}$ has been replaced by the level-$p$ conditional quantile estimation.
In the case one wants to optimize the CSI score, as advocated in Jolliffe (2004), another a more efficient predictor can be built since in this case the optimal probability threshold in Eq. (14) is no longer $p$ as for PSS but is $p^{\prime}$ such that $p^{\prime}=\frac{CSI}{1+CSI}$, which corresponds to $p^{\prime}\approx 0.33$ for an expected CSI around $0.5$. This leads to the near optimal predictor for CSI metrics:
${\widehat{I}}^{3}_{s,d,t+h}={\mathcal{H}}\Big(\widehat{Q}^{s}_{t+h_{i}}(0.33)-Q_{C}^{s}(p)\Big)\;.$ (16)
In section 5.2.3, all these 3 predictors are compared for extreme wind threshold in Corsica.
4 Neural network and baseline models
4.1 Model inputs